OpenStudy (anonymous):

a local kennel hs twice as many cats (c) as dogs (d). when full, the kennel has a total of 30 cats and dogs. which system of equations could be used to find the number of cats in the kennel when it is full?

OpenStudy (danielle):

c+d=30 2c=d so c+2c=30 3c=30 c=10 d=20

OpenStudy (anonymous):

Cats are twice the dogs. So Cats =2 times Dogs Total no= 30 Therefore C + D= 30 C= 2D So substituting C in the equation, 2D+D=30 3D=30 D= 10 Ten Dogs C=20 Twenty Cats

OpenStudy (danielle):

sorry Drishya's answer is correct . I didn't focus on the sentence and tool dogs twice the cats